the doctor orders ceclor 45 mg/kg/day PO in 3 divided doses for a pediatric patient who weighs 66 pounds. the medication comes in 75 mL bottles and is labeled Ceclor 125 mg/mL. how many mL would the nurse administer per dose?
Answer #1

ANSWER.

* Calculation of volume(ml) of drug administer per dose.

        Formula

            Volume(ml) = Desired dose(D)/Dose in hand(H)* Quantity(Q).

Here

       Ordered dose = 45 mg/kg/day ( 3 divided doses)

       Weight of the patient = 66 pounds.

Convert pound to kg for calculation

             1 pound = 0.454 kg

     So

          Weight of the patient = 66 * 0.454

                                            = 29.96 kg.

         Desired dose(D) for this patient = 45 * 29.96

                                                             = 1348.2 mg.

                     Dose in hand (H) = 125 mg.

                             Quantity(Q) = 1 ml

As per above formula

        Volume(ml) of drug required = D/H * Q

                                                      = 1348.2 / 125 * 1

                                                      = 10.79

                                                      = 10.8 ml/day (3 divided doses)

      So

           Volume(ml) of drug administer per dose = 10.8 / 3

                                                                           = 3.6 ml